# OnePerc: A Randomness-aware Compiler for Photonic Quantum Computing

### v2

@@ 7.4. Hyper Parameters
The size of renormalized lattices also relies on a suitable choice of MI ratio (defined in Fig. 8). Fig. 11(c) illustrates the renormalized lattice size with different choices of MI ratios. It can be seens that the renormalization size first increases with the MI ratio and then decreases, peaking at around 7 regardless of the number of modules. This is because an excessively low MI ratio leads to a waste of resource with its wide interval space while an overly high MI ratio increases the probability of unable to connect corresponding paths with its restricted routing space in the intervals.
Average Stripe Size
A suitable choice of average node size is important for reshaping, as it determines the target size of a successful 2D lattice. Fig.12 illustrates the success probability of reaching different predetermined lattice sizes, i.e., different choices of average node size. It can be seen that the success probability approaches 1 rapidly as the target lattice becomes more coarse-grained. This sharp transition motivates us to choose the smallest average node size that brings the success probability close to 1.
Figure 12. Effect of choices of average node size.
@@ 8. Conclusion
In this work, we provide in-depth analysis and discussion of the challenges for photonic quantum compilation brought by the probabilistic operations involved in the computing. We propose a randomness-aware compiler to handle these probabilistic operations, demonstrating a concurrent achievement of scalability and efficiency on photonic systems. With this being said, we believe that there is still significant potential for fully exploring the entire optimization space. We hope that our work could attract more effort from the computer architecture and compiler community to explore the advantages of photonic quantum computing and overcome the unique challenges.

### v3

@@ 7.5. Hyper Parameters
The sizes of renormalized lattices rely on a suitable choice of MI ratio (defined in Fig. 10). Fig. 13(c) illustrates the renormalized lattice size with different choices of MI ratios. It can be seen that the renormalization size first increases with the MI ratio and then slightly decreases, peaking at a value around 7. This is because an excessively low MI ratio leads to a waste of resource with its wide interval space, while an overly high MI ratio increases the probability of unable to connect corresponding paths with its restricted routing space in the intervals.
Average Node Size
A suitable choice of average node size is also important, as it determines the target size of a successful 2D renormalization. Fig.16 illustrates the success probability of reaching different predetermined lattice sizes, i.e., different choices of average node size. It can be seen that the success probability approaches 1 rapidly as the target lattice becomes more coarse-grained. This sharp transition motivates us to choose the smallest average node size that brings the success probability close to 1.
Figure 16. Effect of choices of average node size, with RSL size being $200\times 200$.
@@ 8. Conclusion
In this work, we provide in-depth analysis and discussion of the challenges for photonic quantum compilation brought by the probabilistic operations involved in the computing. We propose a randomness-aware compiler to handle these probabilistic operations, demonstrating a concurrent achievement of scalability and efficiency on photonic systems. Nevertheless, we believe that there is still significant potential for fully exploring the optimization space. We hope that our work could attract more effort from the computer architecture and compiler community to explore the advantages of photonic quantum computing and overcome the unique challenges.
9. Acknowledgement
We thank the anonymous reviewers for their constructive feedback and the cloud bank (cloud_credit, ). This work is supported in part by Cisco Research, NSF 2048144 and Robert N.Noyce Trust.
Appendix A Artifact Appendix
A.1. Abstract
The artifact contains source codes of OnePerc and necessary code scripts to reproduce key results (Table 2, 3, Fig. 12,13,14,15,16) and compare with the baselines in our evaluation. The hardware requirement is a regular X86 server. The software dependencies only contain common python packages. As results in Section 7 are averaged over multiple executions, slight deviation is expected in the reproduction.
A.2. Artifact check-list (meta-information)
  • Algorithm: OnePerc contains two major algorithms.
    • The directory Graph_State_Mapping/ is dedicated to the offline passes, comprising several essential components:
        Construct_Test_Circuit.py creates benchmark circuits with a specific number of qubits.
        Graph_State.py transforms the generated quantum circuits into corresponding program graph states.
        Determine_Dependency.py examines the dependency relationships within the entire graph state.
        Mapping_Routing.py maps the entire graph state onto a virtual hardware of a specific size.
    • The directory Renormalization/ is dedicated to the online passes, comprising the following key components:
        Percolate.py simulates probabilistic fusion within a real physical scenario to generate a physical graph state.
        Renormalization.py reshapes the generated physical graph state to the desired shape of the IR graph state obtained in the offline pass.
        Draw_Grid.py executes 2D renormalization within a single resource state layer.
        Check_Connectivity.py verifies the presence of time-like connections defined in the previous offline pass.
  • Output: The output of the compilation process is the reshaped physical graph state identified within the layers of the physical resource state.
  • Run-time environment: Python, Jupyter Notebook.
  • Hardware: Memory size depends on the benchmark size and whether the refresh is enabled (the largest benchmarks without refresh can be processed with 192 GB RAM).
  • Experiments: Compiling the benchmark programs with OnePerc, using OneQ compiler as the baseline.
  • Required disk space (approximately): When selecting the refresh option, only 32GB of disk space is necessary, whereas opting out could necessitate 130GB of disk space.
  • Metrics: Resource state depth and fusion cost.
  • Time needed to complete experiments: The approximate execution time for each benchmark ranges from 10 seconds to 2 hours with program size expanding for OnePerc. For OneQ basline, the execution time can be infinit. In the experiement setting, it is given a upper bound to cost $10^{6}$ resource state layers, after which the execution time varies from 3 minutes to 6 hours with program size expanding. It will take hundreds of CPU hours to fully reproduce all results in Table 2, 3 and Fig. 12, 13, 14, 15, 16.
  • Publicly available: Yes
  • Code licenses: Apache License 2.0
  • Workflow framework used: Jupyter, Qiskit, PyZX
  • Archived repo: https://zenodo.org/records/10799879
  • DOI: 10.5281/zenodo.10799879
A.3. Description
A.3.1. How to access
This artifact can be downloaded at the link https://zenodo.org/records/10799879.
A.3.2. Hardware dependencies
A standard server with Intel CPUs can effectively run our artifact, with the capacity of RAM potentially constraining the scale of benchmarks that can be executed. In our experiments, we allocated 192GB of RAM to accommodate the execution of all benchmarks. However, activating the refresh option would reduce this requirement to just 32GB of RAM.
A.3.3. Software dependencies
The artifact is developed using Python 3.10, and we require Jupyter Notebook for its utilization. We have prepared files containing scripts to facilitate the automatic and interactive reproduction of results for convenient validation. Pyzx is employed for generating specific quantum circuits, while other dependencies such as NetworkX, Matplotlib, and NumPy are also utilized.
A.4. Installation
To use our artifact, you may download the repo to your local machine from https://zenodo.org/records/10799879 and install the software dependencies by running commands:
conda create -n oneperc python=3.10
pip3 install -r requirements.txt
A.5. Evaluation and expected results
After downloading the artifact and installing all software dependencies, you can open the following jupyter notebook files to reproduce experimental data of baseline and OnePerc for corresponding table and figures.
  • Compiler.ipynb (Table 2)
  • refesh.ipynb (Table 3)
  • sensitivity.ipynb (Fig. 12)
  • scalability.ipynb (Fig. 13, 16)
  • time.ipynb (Fig. 14, 15)
The previous experimental data has already been saved to data/. In scalability.ipynb, sensitivity.ipynb, time.ipynb and refresh.ipynb, setting ‘RunAgain = False’ will generate the plots directly from original data, while setting ‘RunAgain = True’ will run the experiments again, generating new data and new plots. Note that running with a parameter $N$ corresponds to an $N^{2}$-qubit benchmark instead of $N$-qubit.
The generation of Table 2 is the most time-consuming procedure in the evaluation. This is because OneQ performs badly for large programs. Although we force the compilation to terminate when the consumed #RSL reaches $10^{6}$, it can take hours to reach this limit. As a result, we provide three code blocks in Compiler.ipynb.
  • The first code block allows users to run OneQ for individual benchmarks. By changing the value of $N$, users can obtain the result of OneQ for an $N^{2}$-qubit benchmark. We recommend users to try benchmarks from small to large and feel free to stop at the scale they obtain a $10^{6}$RSL for OneQ since larger scales should also lead to a $10^{6}$RSL.
  • The second code block allows users to run OnePerc for individual benchmarks. By changing the value of $N$, users can obtain the result of OnePerc for an $N^{2}$-qubit benchmark. In this process, users can monitor the consumed RAM manually in the terminal.
  • The third code block allows users to obtain all results of OnePerc in Table 2 in one shot.
The experiment results of these code blocks will be automatically saved to data/. After the experiments, users can run Compiler_Table.ipynb to read the saved data and generate Table 2.
Table 3 can be generated by refresh.ipynb, which runs the offline mapping and obtain an estimated #RSL from the number of logical layer. This is because from Fig. 13(b), we know that #RSL has a stable relation with the number of logical layers. The results of non-refreshed #RSL for 25-qubit benchmarks can be obtained directly from Table 2. The ‘-’ in Table 3 means that the compilation utilizes more than 32 GB RAM. In the execution of OnePerc for individual benchmarks (code block 2 in Compiler.ipynb), users can monitor the consumed RAM manually in the terminal (e.g., using htop on Linux). When running the compiler with large enough RAM, users will observe that the consumed RAM exceeds 32 GB for benchmarks larger than 25 qubits. When running the compiler with only 32 GB RAM, users will observe that the compilation of benchmarks larger than 25 qubits would be killed after some time.